the total distance travelled by a boat 1 hour upstream and 30 minutes donwstream in which the rate of current is 3kph
Found 2 solutions by ikleyn, ankor@dixie-net.com:
Answer by ikleyn(52769)

.
It is commonly accepted to print your question, when you post to this forum.

It is A NECESSARY condition to get an answer.


Answer by ankor@dixie-net.com(22740)

the total distance travelled by a boat 1 hour upstream and 30 minutes downstream in which the rate of current is 3kph
:
Let s = the boat speed in still water
then
(s-3) = effective speed upstream
and
(s+3) = effective speed downstream
:
Change 30 min to .5 hr, write a dist equation, dist = time * speed
dist up = dist back
1(s-3) = .5(s+3)
s - 3 = .5s + 1.5
s - .5s = 1.5 + 3
.5s = 4.5
Multiply both sides by 2
s = 9 k/h is the speed in still water
Find the one way distance
1(9-3) = 6 km
Find the total dist up and back: 2 * 6 = 12 km